OpenStudy (anonymous):

Write an equation in slope-intercept form that is perpendicular to 4x – 5y = 1 and passes through (-3, 4). Y=mx+b -5y=-4x+1 -y=-4/5x+1 y=4/5x-1 4=4/5(-3)+b 4=-2.4—-6.4 4=-2.4+6.4 4=4 b=6.4 m=4/5 y=4/5x+6.4 did I do this right?

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (anonymous):

@texaschic101 @Luigi0210

OpenStudy (anonymous):

@jdoe0001 ?

OpenStudy (anonymous):

someone?anyone???????????????????????

OpenStudy (jdoe0001):

$$ 4x-5y=1 \implies -5y=-4x+1 \implies y = \cfrac{-4x+1}{-5}\\ \implies y = \cfrac{-4x}{-5}+\cfrac{1}{-5}\\\\ $$

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!

OpenStudy (jdoe0001):

so, that's the slope for that one, which is PERPENDICULAR to the other one, so, the other will have a slope which is NEGATIVE RECIPROCAL to that, so pretty much just if you have a slope of say "x" x reciprocal of that 1/x negativize it -1/x negative reciprocal

OpenStudy (texaschic101):

4x - 5y = 1 put it in slope intercept form : y = mx +b -5y = -4x + 1 y = (-4x + 1)/-5 y = 4/5x - 1/5 slope is -5/4 (negative reciprocal for perpendicular lines. y = -5/4x + b (sub you points in (-3,4) 4 = - 5/4(-3) + b 4 = 15/4 + b 4 - 15/4 = b 16/4 - 15/4 = b 1/4 = b your equation would be : y = -5/4x + 1/4